OpenStudy (anonymous):

Find the position function s(t) from the given velocity or acceleration function and initial value(s). Assume that units are feet and seconds. v(t) = 40 – sin t, s(0) = 2

OpenStudy (tkhunny):

Position Function: \(x(t) = \int\limits_{0}^{t}v(r)\;dr\) Go!

OpenStudy (anonymous):

I got -1 is that correct?

OpenStudy (tkhunny):

No, it should have a 't' in it. What did you get for \(\int 40 - \sin(t)\;dt\)

OpenStudy (anonymous):

I got (40 - sin(t)) (x)

OpenStudy (tkhunny):

dt not dx.

OpenStudy (tkhunny):

Maybe \(x(t) = 40t + \cos(t) + C\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then i substitute 2 right?

OpenStudy (tkhunny):

Sorry, I've been using x(t). I see the problem statement has specified s(t). \(s(t) = 40t + \cos(t) + C\) No, substitute t = 0, so that s(0) = 2. This establishes the values of "C".

OpenStudy (anonymous):

ohh okay got you